# More than Meets the Eye: Understanding the Effect of Individual Objects on Perceived Visual Privacy

### v1

@@ 3.1.3. Image Generation
(b) Café: Face + Medicine.
(c) Café: Face + Medicine + Landmark.
@@
(e) Office: Face + Tattoo.
(f) Office: Face + Tattoo + Face (a.).
Figure 2. Illustrative examples of image generation in café (top) and office (bottom) environments. Each sequence progresses from a face only foreground PSO to combinations with additional background PSOs.
Figure 2 illustrates examples from both environments. The upper row shows the progression from the base image (generated using café environments) with the face foreground PSO only to variations where medicine and then landmark are added as background PSOs. Similarly, the lower row begins with the base image (generated using office environments), followed by images where the tattoo and then an additional face (a.) are introduced as background PSOs. The complete set of text prompts used to generate these images is provided in Appendix A.1.
@@
3.1.4. Branch Logic Design
@@
Requiring participants to evaluate 48 images could lead to participant fatigue, reduced engagement, and unreliable responses due to rushed decision making. To mitigate this, we used branch logic to direct participants to different flows based on their responses to ranking comfortability levels when the image contains background PSOs. Instead of rating all possible combinations, participants first ranked the five background PSOs by sensitivity. Only the combinations involving the background PSO with the lowest effect on the foreground PSO’s perceived privacy level (PPL) level were then displayed. As a result, participants rated $1+5+4=10$ images per foreground PSO and $30$ images in total. This approach also avoided redundant evaluation: If the PPL of a foreground PSO is highest when paired with a certain background PSO, then requesting participants to reassess it alongside other background PSOs becomes redundant, since the foreground PSO is already perceived as the most sensitive across all pairings. Appendix A.2 shows how the branch logic is implemented when the environment is café, the foreground PSO is face, and the landmark is chosen by the background PSO with the lowest effect.
3.2. Study Practicalities
3.2.1. Survey Organization
@@
The first page of the survey contains the informed consent form, explaining the general objective of the study, how the data will be collected and processed, and compliance with GDPR. Proceeding the survey requires the confirmation of all mandatory consent statements, and an optional checkbox allows entry into a lottery for a movie ticket. Once consent is obtained, participants respond to demographic questions and questions regarding their social media habits and privacy awareness. We constructed privacy awareness questions to understand individual privacy preferences in general. Following this, participants complete three sections, each focusing on a different foreground PSO. In total, the survey contains 50 questions, all of which require a response to proceed, except for the optional open-text fields. We should note that we do not evaluate whether individuals with stronger privacy preferences are more likely to perceive PSOs as private (i.e., less comfortable sharing them online). Our focus is on the effect of context rather than individual privacy awareness. Moreover, prior work (Hoyle et al., 2020) has shown that higher privacy awareness is often associated with perceiving visual content, and attributes shape information sharing behaviors (Bélanger and Crossler, 2011). Nevertheless, we acknowledge that the reported comfortability levels may vary depending on the privacy awareness of the participants. Individuals with stronger privacy concerns might rate PSOs as less comfortable to share, while those with lower awareness may perceive the same objects as less sensitive.
@@
In our user study, each individual section starts with the image containing a single PSO: the foreground PSO, and participants are asked about their comfortability level of sharing this object. The initial question is necessary to obtain a baseline perceived privacy level (PPL). The participants then view a row of images showing the same foreground PSO, paired with a single background PSO different for each image, and a ranking question that asks for the change in their comfortability. They are also asked about the rationale behind their ranking choice (optional open-text answer). The branch logic design is utilized according to their ranking. In the final stage, participants view a row of images including the same foreground PSO, the background PSO ranked with the lowest decrease in their comfortability level (lowest PPL effect) paired with the remaining background PSOs. The final stage also includes an optional question asking to describe whether and in what way the lowest-ranked background PSO influenced the comfortability level of the foreground PSO when combined with other background PSOs.
@@
All ratings were provided on a Likert scale (1-5) where 1 indicated ”not comfortable at all” and 5 ”very comfortable.” We deliberately asked for the comfortability level instead of PPL since it is easier for participants to interpret while still capturing the level of privacy in a more natural form. Table 10 in Appendix A.3 shows the inverse mapping between comfortability level and PPL. The complete set of survey questions is provided in the Appendix A.2.
3.2.2. Survey Deployment
We analyzed commercial platforms in terms of data storage policies, customization levels, and the features they offer, such as complex skip logic scenarios. Then, we selected Streamlit  [^6] https://streamlit.io/ to implement the survey from scratch as it gives the developer full control. We deployed the survey online, and allowed participants to access the survey through a shared link. After completing the survey, the participants submit their responses electronically, and their answers are saved locally without sharing them with any other cloud platform.
3.2.3. Participant Recruitment and Demographics
@@
Participants were recruited using a chain-referral (snowball) sampling approach. We initially distributed the survey link through our personal contacts and social media channels, reaching 73 individuals, and encouraged them to share it further within their own networks. In total, 109 participants completed the survey, with 61 assigned to the café version and 48 to the office version. After excluding responses that showed inconsistencies between ratings and rankings, 92 valid responses remained (51 café and 41 office). As compensation, participants who opted to participate in a raffle were eligible to win one of the 24 movie tickets.
@@
Table 2 shows the demographic distribution of participants. The participants were balanced in sex (51% female, 47% male) and predominantly White/Caucasian (63%) or Asian (22%). Most of the participants’ age was in the range 253425-34 (35%) or 455445-54 (32%), with smaller groups in the ranges 182418-24 (20%), 354435-44 (10%) and $55+$ (4%). Regarding education, the majority had a bachelor’s degree (35%) or a master’s degree (32%), while a quarter (26%) had a doctorate. More than half of the participants were employed full-time (57%), and students made up 30%. Social media usage varied: 37% reported sometimes, 36% rarely, and only 2% daily usage.
Table 2. Demographic distribution of survey participants (n=92).
Category Response Count (%) Category Response Count (%)
Ethnicity White/Caucasian 58 (63%) Education High school graduate 4 (4%)
Asian 20 (22%) Some college 3 (3%)
Black/African 6 (7%) Bachelor’s degree 32 (35%)
Mixed/Multi 3 (3%) Master’s degree 29 (32%)
I don’t wish to disclose 5 (5%) Doctorate 24 (26%)
Age 18–24 18 (20%) Profession Employed full-time 52 (57%)
25–34 32 (35%) Employed part-time 8 (9%)
35–44 9 (10%) Student 28 (30%)
45–54 29 (32%) Retired 2 (2%)
55+ 4 (4%) Unemployed 2 (2%)
Gender Female 47 (51%)
Male 43 (47%)
Other 2 (2%)
3.3. Analysis overview
3.3.1. Quantitative Analysis
To evaluate H1, we used the Wilcoxon signed-rank test as our design compares paired measurements of the same foreground PSO with and without background PSO. The null hypothesis  H1 is “There is no difference between perceived privacy level (PPL) of a foreground PSO when it appears alone versus when it is accompanied by a background PSO”.
To test H2, we compared the comfortability scores between different environments using the Mann-Whitney U test, as different participants rate the same foreground PSO in the café vs. office and the variable of comfortability level is ordinal. The null hypothesis for H2 is “The distributions of perceived privacy levels are the same for café and office”.
For H3, we did not perform statistical testing, as the subset of participants who answered each combination question varied according to their individual rankings, making group-level inference unreliable. Instead, we measured the mean change in comfortability level and the proportion of participants who change their comfortability level when there are two background PSOs.
In addition to the above tests, we used the Mann-Whitney U test (with single categories) to understand whether demographics influence perceived privacy levels. To test factors with more than one category, including age group, profession, and educational level, we used the Kruskal–Wallis H test.
3.3.2. Qualitative Analysis
@@
We conducted qualitative analysis of the open-text responses from the survey questionnaire based on Braun and Clarke’s framework for inductive thematic coding (Braun and Clarke, 2006; Braun et al., 2023). First, the author with an HCI background performed an independent review of the responses to get acquainted with the data using memoing and to generate the initial list of in-vivo codes (Blandford et al., 2016). Then, the author iteratively reviewed the initial codes along with the memos to deduce themes that could represent users’ cognition or behaviour while dealing with visual privacy. This process was repeated until saturation was reached, with no further codes and themes or sub-themes emerging. Finally, the author drafted a codebook along with the definitions, inclusion, and exclusion criteria. The codebook was then reviewed to refine the wording of codes, their definitions, and thematic categories. The final codebook was reviewed once again to ensure that it resonated with the research questions and represented the survey data. Two authors then used the codebook to code all individual responses independently. Since the participants provided their responses as a summary of their cognitive walkthrough during our study, it is possible to tag the responses with multiple codes. Due to this overlap, the coders were allowed to use up to three codes per response. To assess reliability at the theme level, we collapsed the codes into their respective themes. We quantified agreement among coders using Krippendorff’s alpha as the inter-rating reliability (IRR) measure due to its suitability  (McDonald et al., 2019). We found the IRR to be 0.891, which indicates that coders consistently identified the same overarching themes and a strong agreement at the theme level, supporting the reliability of our thematic analysis.
3.4. Study Ethics
This study was conducted in accordance with regional privacy compliance and well-established best practice guidelines for user studies (Redmiles et al., 2017; Rea and Parker, 2014; Egelman et al., 2007). Before the actual survey began, the participants were presented with information about the use of study results and the participation reward terms. Also, the participation was voluntary, and we sought explicit consent from each participant who attempted the survey. The consent and privacy policy included in our study was drafted and provided by the data privacy regulation unit of our institution. A high-level privacy impact assessment was also conducted to review the study methodology and data to be collected to ensure ethical treatment of sensitive data. We ensured that personally identifiable information was solely used for contacting the participants and excluded from the analysis. We used an open-source, GDPR-compliant tool to build and host our study. All the data is entirely under our control on a server in the EU region, and it will be discarded after the project finishes. The visual images used in our study were purely illustrative and did not depict any real human subjects.
4. Results
Before analyzing the results, the responses that showed clear inconsistencies between the rankings and ratings were excluded from the data set. Inconsistency refers to instances where a participant’s assessment of the comfortability level concerning a foreground privacy-sensitive object (PSO), when there is a single background PSO, conflicts with the subsequent ranking question. This method also enabled us to filter out inattentive participants, thereby improving the quality of the remaining data.
4.1. Effect of Users’ Privacy Attitudes
Table 3 presents the participants’ responses to questions concerning privacy awareness. Concern about online privacy (Q1, 90% agreed) and discomfort with the disclosure of personal information to online companies (Q3, 94% agreed) received the highest ratings. Concern about privacy in everyday life (Q2, 78% agreed) was also significant but slightly lower overall. From the responses, we inferred that the majority of the participants expressed their concerns about their privacy online, in everyday life, and with the companies. Responses to (Q4) show that the participants’ ratings of how often they check for personal information before sharing a photo on social media were lower than their ratings of concern for privacy online. This indicates that while most participants express concern about online privacy, this concern does not consistently translate into proactive behaviors, including looking for personal information before sharing it on social media, highlighting a gap between privacy attitudes and practices.
@@ 4.1. Effect of Users’ Privacy Attitudes
Table 3. Responses to Privacy Awareness Questions (%) and Mean Value. (SD = Strongly Disagree (1), SWD = Somewhat Disagree (2), N = Neutral (3), SWA = Somewhat Agree (4), SA = Strongly Agree (5)).
@@
4.2. Effect of Background PSOs (H1)
To examine H1, we quantitatively compared comfortability levels across environments. As shown in Figure 3, the mean ratings (averaged across café and office environments) illustrate how the presence of background PSOs influences the comfortability level (the perceived privacy level, PPL) of each foreground PSO. In each figure, the first bar indicates the baseline comfortability level when the image contains only the foreground PSO, while the subsequent bars represent its combination with different background PSOs. For Face, the strongest reduction occurs with medicine (mean = 1.65; –1.48 from the baseline of 3.13), followed by the presence of face (a.) (2.61; –0.52). In contrast, tattoo (2.99) and landmark (2.96) produce minor decreases. For Miscellaneous Paper, full name shows the smallest drop (2.12; 0.28-0.28), whereas address (1.39) and signature (1.38) yield the largest decreases. For Computer/phone Screen, identifiers such as email (1.62), username (1.43), and phone number (1.48) lead to strong reductions, face (r.) causes a moderate decrease (2.21). Interestingly date stands out as the only case associated with an increase. A possible reason for this might be the averaging of rankings from both settings, which influences the participants’ judgements. Overall, the addition of background PSOs nearly always elevates the PPL of the foreground PSO, with certain categories (e.g., medicine for face, signature for miscellaneous paper, username for screen) exerting the most pronounced effects.
@@ 4.2. Effect of Background PSOs (H1)
Figure 3. Average comfortability level ratings for foreground PSOs and the effect of a single background PSO.
Figure 3 shows only mean values and does not capture the underlying distributions. To systematically examine whether observed differences were consistent, we compared the comfortability levels of each foreground PSO alone with their corresponding foreground/background PSO pair using the Wilcoxon signed-rank test (see Section 3.3.1). Table 4 summarizes the results. Overall, the presence of background PSOs often reduced the comfortability level, indicating an increase in perceived privacy level (PPL).
For the Face, the addition of face (a.) significantly increased PPL in the office but not in the café, suggesting that the effect of background PSO matters more in professional contexts. Medicine strongly increased PPL in both environments, while tattoo and landmark had no significant effect.
For Miscellaneous Paper, address, birth date and signature consistently raised PPL across both environments. Face (p) increased PPL only in office, while full name had no significant effect.
For Computer/phone Screen, nearly all background PSOs increased PPL, except date, which has no effect.
Our analysis shows that not all background PSOs equally affect the perceived privacy. Highly sensitive cues (e.g., medicine, addresses, signatures, digital identifiers on screens) consistently amplify the PPL of foreground PSOs, while more ambiguous cues (e.g., tattoos, landmarks, dates) often ignored. This pattern reflects with everyday social media behaviour, where users avoid sharing explicit sensitive information such as medicine or ID cards, but tend to underestimate the risks posed by ambigious cues that can be exploited in inference attacks (e.g., social engineering, location tacking, identity linking, etc.).
@@
Table 4. Wilcoxon signed-rank test results: Change in comfortability level ($\Delta$ mean) when adding a background PSO. Significant $p$-values (<0.05<0.05) are highlighted in bold.
Foreground Background Café ($\Delta$, $p$) Office ($\Delta$, $p$) Significance
Face Face (a.) -0.37 (0.0576) -0.71 (0.0003) Only Office
Poster -0.55 (0.0345) -0.22 (0.2193) Only Café
Medicine -1.78 (0.0000) -1.10 (0.0000) Both
Tattoo -0.22 (0.4359) -0.05 (0.6917) None
Landmark -0.16 (0.4115) -0.20 (0.1895) None
Misc. Paper Face (p.) -0.33 (0.0815) -0.59 (0.0049) Only Office
Full Name -0.25 (0.1256) -0.32 (0.0920) None
Address -1.22 (0.0000) -0.76 (0.0006) Both
Birth Date -0.57 (0.0039) -0.73 (0.0012) Both
Signature -1.16 (0.0000) -0.85 (0.0001) Both
Computer/phone Face (r.) -0.47 (0.0165) -0.80 (0.0001) Both
Screen Date +0.25 (0.0849) +0.02 (0.7181) None
E-mail -1.37 (0.0000) -1.00 (0.0005) Both
Username -1.57 (0.0000) -1.17 (0.0000) Both
Phone Number -1.59 (0.0000) -1.05 (0.0001) Both
4.3. Effect of Environment (H2)

### v2

@@ III-B3 Image Generation
(b) Face + Medicine.
(c) Face + Medicine + Landmark.
@@
(e) Face + Tattoo.
(f) Face + Tattoo + Face (a.).
Figure 2: Illustrative examples of image generation in café (top) and office (bottom) environments. Each sequence progresses from a face only foreground PSO to combinations with additional background PSOs.
@@
III-B4 Branch Logic Design
@@
Requiring participants to evaluate 48 images could lead to participant fatigue, reduced engagement, and unreliable responses due to rushed decision making. To mitigate this, we used branch logic to direct participants to different flows based on their responses to ranking comfortability levels when the image contains background PSOs. Instead of rating all possible combinations, participants first ranked the five background PSOs by sensitivity. Only the combinations involving the background PSO with the lowest effect on the foreground PSO’s perceived privacy level (PPL) level were then displayed. As a result, participants rated $1+5+4=10$ images per foreground PSO and $30$ images in total. Each participant views identical $3\times(1+5)=18$ images from the same environment, while the last 12 differ due to branching. This approach also avoided redundant evaluation: If the PPL of a foreground PSO is highest when paired with a certain background PSO, then requesting participants to reassess it alongside other background PSOs becomes redundant, since the foreground PSO is already perceived as the most sensitive across all pairings. Appendix A-B shows how the branch logic is implemented when the environment is café, the foreground PSO is face, and the landmark is chosen by the background PSO with the lowest effect.
III-C Study Practicalities
III-C1 Survey Organization
@@
The first page of the survey contains the informed consent form, explaining the general objective of the study, how the data will be collected and processed, and compliance with GDPR. Proceeding the survey requires the confirmation of all mandatory consent statements, and an optional checkbox allows entry into a lottery for a movie ticket. Once consent is obtained, participants respond to demographic questions and questions about their social media habits and privacy awareness. The privacy awareness questions included four brief items to understand individual privacy preferences in general. Following the approach of Hoyle et al. [hoyle2020privacynorms], we opted for a short scale with minimal burden on respondents rather than a full psychometric scale such as the IUPUC [malhotra2004internet]. Following this, participants complete three sections, each focusing on a different foreground PSO. In total, the survey contains 50 mandatory questions, except for the optional open-text fields. We should note that we do not evaluate whether individuals with stronger privacy preferences are more likely to perceive PSOs as private (i.e., less comfortable sharing them online). Our focus is on the effect of context rather than individual privacy awareness. Moreover, previous work [hoyle2020privacynorms] has shown that higher privacy awareness is often associated with perceiving visual content, and attributes shape information sharing behaviors [Belanger2016Privacy]. Nevertheless, we acknowledge that the reported comfortability levels may vary depending on the privacy awareness of the participants. Individuals with stronger privacy concerns might rate PSOs as less comfortable to share, while those with lower awareness may perceive the same objects as less sensitive.
@@
In our user study, each section begins with an image containing a single foreground PSO, and participants are asked about their comfortability level of sharing this object. The initial question is necessary to obtain a baseline perceived privacy level (PPL). The participants then view a row of images showing the same foreground PSO that is paired with a single background PSO different for each image, and are asked for the comfortability level for the same foreground PSO. They also rank the effect of background PSO on the degree of change and are asked about the rationale behind their ranking choice (optional open-text answer). The branch logic design is utilized according to their ranking. In the final stage, participants view a row of images including the same foreground PSOs, the background PSO ranked with the lowest decrease in their comfortability level (lowest PPL effect) paired with the remaining background PSOs. The final stage also includes an optional question asking to describe whether and in what way the lowest-ranked background PSOs influenced the comfortability level of the foreground PSO when combined with other background PSOs.
@@
All ratings used a 5-point Likert scale, where 1 indicated “not comfortable at all” and 5 “very comfortable”. We asked for comfortability level instead of PPL since it is easier for participants to interpret while still naturally capturing the perceived privacy. Table IX in Appendix A-C shows the inverse mapping between the comfortability level and PPL. The complete survey questionnaire is provided in the Appendix A-B.
III-C2 Survey Deployment
Based on a careful review of the data storage policies, customizability, and control logic offered by different survey platforms, we selected Streamlit  [^3] https://streamlit.io/ to implement the survey from scratch. We deployed the survey online and provided participants with a shared link to access it. Streamlit stores responses locally during completion and sends them to a designated email address upon submission. We downloaded these responses to a local machine for analysis.
III-C3 Participant Recruitment and Demographics
@@
Participants were recruited using a chain-referral (snowball) sampling approach. We initially distributed the survey link through our personal contacts and social media channels, reaching 73 individuals, and encouraged them to share it further within their own networks. In total, 109 participants completed the survey, with 61 assigned to the café version and 48 to the office version. After excluding responses that showed inconsistencies between ratings and rankings, 92 valid responses remained (51 café and 41 office). Participants were compensated with movie tickets through voluntary participation in a raffle.
@@
Table X shows the demographic distribution of the participants. The participants were balanced in sex (51% female, 47% male) and predominantly White/Caucasian (63%) or Asian (22%). Most of the participants’ age was in the range 253425-34 (35%) or 455445-54 (32%), with smaller groups in the ranges 182418-24 (20%), 354435-44 (10%) and $55+$ (4%). Regarding education, the majority had a bachelor’s degree (35%) or a master’s degree (32%), while a quarter (26%) had a doctorate. More than half of the participants were employed full-time (57%), and students made up 30%. Social media usage varied: 37% reported sometimes, 36% rarely, and only 2% daily usage.
III-D Analysis overview
III-D1 Quantitative Analysis
To evaluate H1, we used the non-parametric Wilcoxon signed-rank test, since our design compares paired measurements of the same foreground PSO with and without background PSO (repeated measure with post intervention). The null hypothesis  H1 is “There is no difference between the perceived privacy level of a foreground PSO when it appears alone versus when accompanied by a background PSO”. To test H2, we compared the comfortability scores between different environments using the non-parametric Mann-Whitney U test, as each participant rate the same foreground PSO only in one of the environments (café vs. office as independent variable). The null hypothesis for H2 is “The distributions of perceived privacy levels are the same for café and office”. For H3, we did not perform statistical testing, as the subset of participants who answered each combination question varied by individual rankings, making group-level inference unreliable (less number of repeated measures). Instead, we report the mean change in comfortability level and the proportion of participants who changed their comfortability level when two background PSOs were present. In addition these tests, we used the Mann-Whitney U test (with single categories) to understand whether demographics influence perceived privacy levels. To test factors with more than two categories, including age group, profession, and educational level, we used the Kruskal–Wallis H test.
III-D2 Qualitative Analysis
@@
We conducted a qualitative analysis of the open-text responses from the survey questionnaire using inductive, in-vivo coding followed by thematic clustering to group related codes into higher-order themes [braun2023doing, blandford2016qualitative]. First, the author with an HCI background performed an independent review of the responses to get acquainted with the data using memoing and to generate the initial list of codes. Then, the author iteratively reviewed the initial codes along with the memos to deduce themes that could represent users’ cognition or behaviour while dealing with visual privacy. This process was repeated until saturation was reached, with no further codes and themes or sub-themes emerging. Finally, the author drafted a codebook along with the definitions, inclusion, and exclusion criteria. The codebook was then reviewed to refine the wording of codes, their definitions, and thematic categories. The final codebook was reviewed once again to ensure that it resonated with the research questions and represented the survey data. Two authors then used the codebook to code all individual responses independently. Since the participants provided their responses as a summary of their cognitive walkthrough during our study, it is possible to tag the responses with multiple codes. Due to this overlap, the coders were allowed to use up to three codes per response. To assess reliability at the theme level, we collapsed the codes into their respective themes. We quantified agreement among coders using Krippendorff’s alpha as the inter-rating reliability (IRR) measure due to its suitability  [mcdonald2019reliability]. We found the IRR to be 0.891, which indicates that coders consistently identified the same overarching themes and a strong agreement at the theme level, supporting the reliability of our thematic analysis.
@@
TABLE II: Responses to Privacy Awareness Questions (%) and Mean Value. (SD = Strongly Disagree (1), SWD = Somewhat Disagree (2), N = Neutral (3), SWA = Somewhat Agree (4), SA = Strongly Agree (5)).
@@
III-E Study Ethics
This study was conducted in accordance with regional privacy compliance and well-established best practice guidelines for user studies [redmiles2017summary, rea2014designing, egelman2007security]. Before the actual survey began, the participants were presented with information about the use of study results and the participation reward terms. The participation was voluntary, and we sought explicit consent from each participant who attempted the survey. The consent and privacy policy included in our study was drafted and provided by the data privacy regulation unit of our institution. A high-level privacy impact assessment was also conducted to review the study methodology and data to be collected to ensure ethical treatment of sensitive data. We ensured that personally identifiable information was solely used for contacting the participants and excluded from the analysis. We used an open-source, GDPR-compliant tool to build and host our study. All the data is entirely under our control on a server in the EU region, and it will be discarded upon the camera-ready version of the manuscript. The visual images used in our study were purely illustrative and did not depict any real human subjects.
IV Results
Before analyzing the results, the responses that showed clear inconsistencies between the rankings and ratings were excluded from the dataset. Inconsistency refers to instances where a participant’s assessment of the comfortability level concerning a foreground PSO, when there is a single background PSO, conflicts with the subsequent ranking question. This method also allowed us to filter out inattentive participants, thereby improving the quality of the remaining data.
IV-A Effect of Users’ Privacy Attitudes
Table II presents the participants’ responses to questions concerning privacy awareness. Concern about online privacy (Q1, 90% agreed) and discomfort with the disclosure of personal information to online companies (Q3, 94% agreed) received the highest ratings. Concern about privacy in daily life (Q2, 78% agreed) was also significant but slightly lower overall. From the responses, we inferred that the majority of the participants expressed their concerns about their privacy online, in everyday life, and with the companies. The responses to (Q4) show that the participants’ ratings of how often they check personal information before sharing a photo on social media were lower than their ratings of concern for privacy online. This indicates that while most participants express concern about online privacy, this concern does not consistently translate into proactive behaviors, including looking for personal information before sharing it on social media, highlighting a gap between privacy attitudes and practices.
IV-B Effect of Background PSOs (H1)
To examine H1, we quantitatively compared comfortability levels and averaged using both environments. Figure 3 illustrates how the presence of background PSO influences the comfortability level (the perceived privacy level, PPL) of each foreground PSO. In each figure, the first bar indicates the baseline comfortability level when the image contains only the foreground PSO, while the subsequent bars represent its combination with different background PSOs. For Face, the strongest reduction occurs with medicine (mean = 1.65; –1.48 from the baseline of 3.13), followed by face (a.) (2.61; –0.52). In contrast, tattoo (2.99) and landmark (2.96) produce minor decreases. For Miscellaneous Paper, full name shows the smallest drop (2.12; 0.28-0.28), whereas address (1.39) and signature (1.38) yield the largest decreases. For Computer/phone Screen, identifiers such as email (1.62), username (1.43), and phone number (1.48) lead to strong reductions, face (r.) causes a moderate decrease (2.21). Interestingly date stands out as the only case associated with an increase. A possible reason for this might be the averaging of rankings from both settings, which influences the participants’ judgements. Overall, the addition of background PSOs mostly elevates the PPL of the foreground PSO, with certain categories (e.g., medicine for face, signature for miscellaneous paper, username for screen) exerting the most pronounced effects.
@@ IV-B Effect of Background PSOs (H1)
Figure 3: Average comfortability level ratings for foreground PSOs and the effect of a single background PSO.
@@
TABLE III: Wilcoxon signed-rank test results: Change in comfortability level ($\Delta$ mean) when adding a background PSO. Significant $p$-values (<0.05<0.05) are highlighted in bold and blue.
Foreground PSO Background PSO Café ($\Delta$, $p$) Office ($\Delta$, $p$) Significance Face Face (a.) -0.37 (0.0576) -0.71 (0.0003) Only Office Poster -0.55 (0.0345) -0.22 (0.2193) Only Café Medicine -1.78 (0.0000) -1.10 (0.0000) Both Tattoo -0.22 (0.4359) -0.05 (0.6917) None Landmark -0.16 (0.4115) -0.20 (0.1895) None Misc. Paper Face (p.) -0.33 (0.0815) -0.59 (0.0049) Only Office Full Name -0.25 (0.1256) -0.32 (0.0920) None Address -1.22 (0.0000) -0.76 (0.0006) Both Birth Date -0.57 (0.0039) -0.73 (0.0012) Both Signature -1.16 (0.0000) -0.85 (0.0001) Both Computer/phone Screen Face (r.) -0.47 (0.0165) -0.80 (0.0001) Both Date +0.25 (0.0849) +0.02 (0.7181) None E-mail -1.37 (0.0000) -1.00 (0.0005) Both Username -1.57 (0.0000) -1.17 (0.0000) Both Phone Number -1.59 (0.0000) -1.05 (0.0001) Both
Figure 3 shows only mean values and does not capture the underlying distributions. To systematically examine whether observed differences were consistent, we compared the comfortability levels of each foreground PSO alone with their corresponding foreground/background PSO pair using the Wilcoxon signed-rank test. Table III shows that the presence of background PSOs often reduced the comfortability level, indicating an increase in perceived privacy level (PPL). For the Face, the addition of face (a.) significantly increased PPL in the office but not in the café, suggesting that the effect of background PSO matters more in professional contexts. Medicine strongly increased PPL in both environments, while tattoo and landmark had no significant effect. For Miscellaneous Paper, address, birth date, and signature consistently raised PPL across both environments. Face (p) increased PPL only in office, while full name had no significant effect. For Computer/phone Screen, nearly all background PSOs increased PPL, except date, which has no effect.
Our analysis shows that not all background PSOs equally affect perceived privacy. Highly sensitive cues (e.g., medicine, addresses, signatures, digital identifiers on screens) consistently amplify the PPL of foreground PSOs, while more ambiguous cues (e.g., tattoos, landmarks, dates) are often ignored. This pattern reflects with everyday social media behavior, where users avoid sharing explicit sensitive information such as medicine or ID cards, but tend to underestimate the risks posed by ambigious cues that can be exploited in inference attacks such as social engineering or identity linking.
IV-C Effect of Environment (H2)